1. ) The following table shows the number of units produced each month and the total incurred.


MONTH UNITS COST

JUNE 200 50,000

JULY 500 75,000

AUGUST 300 55,000

SEPTEMBER 800 95,000

OCTOBER 700 80,000

NOVERMBER 600 80,000

DECEMBER 400 60,000


Required

              I. Estimate the variable cost per unit and the fixed cost per month use high low method. (6mks)


             ii.  Calculate the regression line Y= a+bx (6mks)



(b) Explain four purpose of transfer pricing (8mks)




1
Expert's answer
2022-02-17T18:01:17-0500

Variable Cost Per Unit = (Highest Activity Cost – Lowest Activity Cost) / (Highest Activity Units – Lowest Activity Units)

80095000 -20050000=60045000


31_29=2 60045000/2 =30022500


Fixed cost ,=Fixed Cost = Highest Activity Cost – (Variable Cost Per Units * Highest Activity Units

80095000-(30022500×31) =850602500


2.The equation has the form Y= a + bX, where Y is the dependent variable (that's the variable that goes on the Y axis), X is the independent variable (i.e. it is plotted on the X axis), b is the slope of the line and a is the y-intercept. Y=a+bx

3.1) Maximizing overall after-tax profits. 3) Circumventing the quota restrictions (in value terms) on imports. ADVERTISEMENTS: 4) Reducing exchange exposure, circumventing exchange controls and restricting profit repatriation so that transfer firms affiliates to the parent can be maximized.